# Revisiting Your Memory: Reconstruction of Affect-Contextualized Memory via EEG-guided Audiovisual Generation

### v1

@@ Introduction
Table 1: Quantitative results calculating distances between text prompt and output (thumbnail image and music) at CLIP and CLAP spaces.
Method
Extracting Human Affects
To effectively extract affects from individual subjects within a single session of neural recordings, we utilized the contrastive learning-based neural network, CEBRA (Schneider, Lee, and Mathis 2023). By leveraging temporally synchronized behavioral sequences as auxiliary variables, CEBRA non-linearly reduces the dimensions of neural signals to maximize the separation between distinct behavioral states while clustering signals from similar states. Initially validated in animal neuroscience (Kim et al. 2024) and invasive BCI applications (Merk et al. 2023), CEBRA has recently demonstrated superior performance in human affect decoding during movie viewing (Yi et al. 2024), suggesting its potential for EEG-based affect decoding.
In our study, CEBRA utilizes keypressed valence sequences (neutral, positive, negative) as auxiliary variables to identify individual-level latent clusters for each affective state. We performed multi-session training of CEBRA with 10 participants’ EEG and key-pressed valence sequence, aligning individuals’ latent valence-neural representations. To examine the generalizability of embeddings, we conducted leave-one-out valence decoding tasks. We trained a k-nearest neighbor (KNN) classifier with 9 participants’ CEBRA embeddings to predict valence label at each timepoint and evaluated its predictive performances with the other’s embeddings.
Aligning Extracted Affects With Text
To effectively incorporate extracted affects into audio/visual output, it is essential to convert them into text prompts, which are commonly used for directing generative processes. Specifically, we pre-defined descriptive terms for positive and negative affective states with the pre-trained language model, claude.ai. For each affective state, we randomly selected words from the predetermined word samples and fine-tuned the original text prompts that the subject articulated in their memories with affective samples using the pre-trained language model. In the neutral state, the original text prompts remain unaltered.
Implementation Details
Affect Extraction. To decode each individual’s affect dynamics from memory-recall EEG signals, we conducted multi-session training of CEBRA. We set hyperparameters as follows: batch size=2,048, model architecture=‘offset-10 model’, the number of hidden units=95, learning rate=.005, the number of iterations=2000, hybrid=False. Following the previous study on the optimal number of dimensionality for generalizable valence representation across individuals and stimuli (Yi et al. 2024), we set the number of latent embeddings’ dimensionality as 7. The neighborhood parameters of KNN classifiers were fixed at the square of the number of the input time points.
Music and Video Generation. Using MusicGen-melody (Copet et al. 2024), we generated music using affect-specific text prompts and memory-associated guiding melodies selected by participants. The generated content was then segmented according to affective state durations and integrated into a final sample using 40 ms crossfading to match the affective trajectory. For video generation, we used stable diffusion ver.1.5 (Rombach et al. 2022). In all our experiments, we fix the parameters of text encoders and LDMs for music and video generation, respectively. We use the author-released codes using default configurations. All experiments were conducted using the PyTorch framework (Paszke et al. 2019) on a single NVIDIA A100(40G) GPU.
Experimental Results
Affect Decoding. Participants showed individual differences in temporal dynamics of affective states during memory recall and the duration of the recall (see Figure 3a). Despite these variations, our CEBRA model demonstrated a plateau in learning performance in multiple-session training with EEG signals and real-time valence keypresses from nine participants (see Figure 3b). The individually derived latent valence-neural embeddings from this model revealed a consistent geometric structure across individuals (see Figure 3c). Using these embeddings with a KNN classifier for leave-one-out classification for each individual yielded a test-weighted F1 score of 0.9.
Qualitative Evaluation. Figure 4 illustrates that our approach produces qualitatively satisfactory visual and audio outputs across all affective states and subjects. Our model preserves the elements of the original sketch while integrating corresponding affects through tones, colors, and visual composition. The positive video clip showed brighter colors and a vibrant tone throughout the videos, whereas the negative displayed darker tones and a gloomier atmosphere. Moreover, the generated output demonstrated satisfactory visual coherence across video frames. Also, our model successfully transfers the style of original guiding melodies. As in Figure 4, the positive affects reflecting music clips showed broader frequency distributions with stronger energy in high-frequency regions, creating brighter and more vibrant spectral patterns. In contrast, the negative affects reflecting content showed concentrated energy in lower frequency bands, resulting in darker and more focused spectral patterns. These results demonstrated clear acoustic differences between affective expressions in generated music. The down-sampled version of generated music samples can be accessed within the supplementary materials.
Quantitative Evaluation. We show quantitative evaluation in Tab. 1 and Figure 5. Semantic evaluation using CLIP and CLAP embeddings shows consistent distances between reference prompts and affect-contextualized outputs across affective states, indicating semantic coherence. Additionally, the minimal distances between outputs and their corresponding affective word embeddings (e.g., positive-positive pairs) validate our model’s ability to effectively direct content generation toward intended affective states.
We examined how affective inputs create variations in visual (e.g., hue, saturation, value) and musical (e.g., mode, brightness, intensity) attributes of generated content. While visual generation shows significant correlations between affects and hue/value, musical generation demonstrates stronger correlations across all features, suggesting more effective affective expression through musical elements.
@@
Figure 6: User study results. (a) Cross-correlation coefficients of real-time valence keypresses between the session 1 and 2. (b) preference assessment results.
@@
User Study. Participants showed valence dynamics more closely aligned with their valence reports during Session 1 when viewing the CEBRA-decoded true video compared to the fake video (see Figure 6a). A Wilcoxon rank-sum test on the best cross-correlation coefficients between real-time valence keypresses in Session 1 and 2 revealed that the true condition produced significantly higher coefficients than the perm condition (P = .012). Additionally, about 56% of participants reported that the true video better reflected their memories and associated affective changes (see Figure 6b). In a preference rating, 5 out of 9 participants chose the true video. These results suggest that our approach can be promising to capture one’s idiosyncratic affect dynamics in their personal memories and generate them as music videos.
Conclusion & Future Works
@@
This research makes significant contributions to cognitive/affective neuroscience and brain-computer interfaces (BCI). We introduce RecallAffectiveMemory, a novel task for studying dynamic affective states during autobiographical memory recall, supported by a comprehensive multimodal dataset (EEG-AffectiveMemory). We demonstrate successful temporal affective trajectory decoding using CEBRA and present RYM, a framework generating personalized audiovisual content synchronized with individual affective experiences.
@@
Despite successfully generating affect-contextualized content, future work will address two key limitations: improving the smoothness of transitions between affective states in the generated content and exploring the feasibility of end-to-end generative models for more seamless integration of affective trajectories. Further research will also focus on broader applications, including therapeutic interventions and enhanced human-computer interaction.

### v2

@@ 5.2. Qualitative Evaluation
Table 1. Quantitative results calculating distances between text prompt and output (thumbnail image and music) at CLIP and CLAP spaces.
@@
Figure 6. User study results. (a) Cross-correlation coefficients of real-time valence keypresses between the session 1 and 2. (b) Preference assessment results.
[…]
5.3. Quantitative Evaluation
We show quantitative evaluation in Tab. 1 and Figure 5. Semantic evaluation using CLIP and CLAP embeddings shows consistent distances between reference prompts (i.e., prompts without affective words) and affect-contextualized outputs across affective states, indicating semantic coherence. Additionally, the minimal distances between outputs and their corresponding affective word embeddings (e.g., positive-positive pairs) validate our model’s ability to effectively direct content generation toward intended affective states.
We analyzed whether the generated videos exhibited variations in visual (e.g., hue, saturation, value) and musical features (e.g., mode, emotional brightness, intensity) depending on participants’ affective states. To this end, we computed eta-squared, the effect size of a one-way ANOVA, to assess whether each audio and visual feature showed significant differences across three affective states for each participant. Results revealed that, except for saturation, all audiovisual features showed larger effect sizes in the true condition compared to the fake condition (Figure 5a), indicating that the videos in the true condition more accurately reflected participants’ valence-dependent characteristics. However, within the true condition, we observed considerable individual differences in the effect sizes of affective states across features (Figure 5b). For example, while subject 2’s video displayed large changes in musical features based on affective states, subject 6’s clip showed notable effect sizes of valence in visual features.
5.4. User Study
@@
Participants showed valence dynamics more closely aligned with their valence reports during Session 1 when viewing the CEBRA-decoded true video compared to the fake video (see Figure 6a). A Wilcoxon rank-sum test on the best cross-correlation coefficients between real-time valence keypresses in Session 1 and 2 revealed that the true condition produced significantly higher coefficients than the perm condition (${\mathbf{r}}_{true}=0.265$, ${\mathbf{r}}_{perm}=0.095$, p = .012). Additionally, about 56% of participants reported that the true video better reflected their memories and associated affective changes (see Figure 6b). In a preference rating, 5 out of 9 participants chose the true video. These results suggest that our approach can be promising to capture one’s idiosyncratic affect dynamics in their personal memories and generate them as music videos.
6. Conclusion & Future Works
@@
This research makes significant contributions to cognitive/affective neuroscience and BCI. We introduce RevisitAffectiveMemory, a novel task for studying dynamic affective states during autobiographical memory recall, supported by a comprehensive multimodal dataset (EEG-AffectiveMemory). We demonstrate successful temporal affective trajectory decoding using CEBRA and present RYM, a framework generating personalized audiovisual content synchronized with individual affective experiences.
@@
Our study aimed to provide a proof-of-concept, demonstrating the feasibility of our proposed framework. To this end, we intentionally selected a homogeneous sample (right-handed Korean young adults, ages 22-28, without specific neurological conditions) to control for potential confounding variables such as cultural background, age, and mental wellness. Similarly, we utilized established neural network components for our framework to ensure extensibility and compatibility with various generative models. While these methodological choices focused our assessment on the framework’s capability, they introduced limitations in sample diversity and technical novelty. Nevertheless, our sample size aligns with prior foundational studies in EEG-based visual reconstruction (Shimizu and Srinivasan, 2022; Lee et al., 2021), and our approach successfully reconstructs affect-contextualized memories, highlighting the potential of our novel task for AI-driven personal content generation. In future studies, we will test the broader applicability of our framework with larger and more diverse samples to account for the idiosyncratic nature of emotional memory. Also, a more systematic investigation into the interaction between the characteristics of recalled memories and the features of the generated videos is warranted in future research (e.g., how do generated videos differ between memories of relatively recent events and those from early childhood?). Last, we will improve technical limitations of our framework, such as the smoothness of transitions between affective states in the generated content.
Acknowledgements.
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2021R1C1C1006503, RS-2023-00266787, RS-2023-00265406, RS-2024-00421268, RS-2024-00342301, RS-2024-00435727, NRF-2021M3E5D2A01022515), by Creative-Pioneering Researchers Program through Seoul National University (No. 200-20240057, 200-20240135), by Semi-Supervised Learning Research Grant by SAMSUNG (No.A0342-20220009), by Identify the network of brain preparation steps for concentration Research Grant by LooxidLabs (No.339-20230001), by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)] by the MSIT (Ministry of Science, ICT), Korea, under the Global Research Support Program in the Digital Field program (RS-2024-00421268) supervised by the IITP (Institute for Information & Communications Technology Planning & Evaluation), by the National Supercomputing Center with supercomputing resources including technical support (KSC-2023-CRE-0568), by the Ministry of Education of the Republic of Korea and the National Research Foundation of Korea (NRF-2021S1A3A2A02090597), by the Korea Health Industry Development Institute (KHIDI), and by the Ministry of Health and Welfare, Republic of Korea (HR22C1605), by Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT (MSIT, Korea) & Gwangju Metropolitan City and by KBRI basic research program through Korea Brain Research Institute funded by Ministry of Science and ICT (25-BR-05-01). This work was also supported by the U.S. Department of Energy (DOE), Office of Science (SC), Advanced Scientific Computing Research program under award DE-SC-0012704 and used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award ASCR-ERCAP0033711.